# Iceberg Sensemaking: A Process Model for Critical Data Analysis

### v3

@@ 7.1 Limitations
A fundamental limitation of all interpretivist approaches, including our own, is that centering subjectivity necessarily limits the reproducibility and replicability of results [20]. Two analysts who use Iceberg Sensemaking may well come to different conclusions. Our model locates this difference in the acknowledgement of tacit schemas, where the specific information one chooses to articulate from the mass of submerged tacit schemes may be entirely contingent, leading to different sensemaking outcomes between different analysts in varying conditions and contexts. When reproducible results are required, a broadly positivist approach is still tenable under this model because the epistemic pluralism of Iceberg Sensemaking is inclusive of positivism, as well—just as long as the limitations of a positivist approach have been sufficiently scrutinized through the broader interpretivist framework of this model. Moreover, positivist approaches will typically be more time-efficient than the slower interpretivist approaches that demand scrutiny of the received wisdom and prevailing expectations of a scientific field.
@@ 7.2 The Value of an Interpretivist Model
In building this model, we have made an active effort to describe how positivist approaches to data analysis work when they work, while also explaining how they go wrong. A sensemaking model should be just as good at diagnosing how the process fails as describing how it succeeds. While the positivism of past sensemaking models is often tacit, it is most evident in the absence of consideration for faulty and misleading data. Correcting this limitation requires more than just the inclusion of a cleaning phase, because the deepest problems with many datasets only emerge through scrutiny of conceptual frameworks that are often taken for granted.
@@
As a rule of thumb, the more political a subject, the more fraught the implications of tacit schemas in data analysis, and the greater the need for interpretivist approaches to sensemaking. For example, court records may seem like straightforward data, but they carry a tacit schema reflecting the entire structure of the criminal justice system. Likewise, calibrating a sensor to measure water quality may seem straightforward, but making use of the readings to assess public health concerns involves acts of interpretation that are deeply embedded in social context. While data entry and automated data gathering could be seen as passive tasks that simply require application of the intended cognitive schema, sensemaking with these datasets requires active attention to the construction of interpretive schemas and articulation of the tacit schemas already in play. Although the causes of data bias are typically discussed in the context of ethics, they are just as much an epistemic matter in cases where unfairness is rooted in failures of investigative rigor.
@@
In the specific case of visualization, the nested model of visualization design (Munzner 2009) can be a useful guide for when to take an interpretivist or cognitivist stance. In the nested model, the abstraction layer is where the designer must make judgements about the tasks and data of target users, and this is prime territory for an interpretivist approach. Problem-driven design and requirements elicitation are processes that necessarily involve interpretation, judgment, and subjectivity. In contrast, the lower two levels of the nested model are more amenable to a cognitivist approach. At the idiom layer, which deals with visual encoding and interaction design choices, success is often gauged using methods such as controlled laboratory experiments where human performance is measured in terms of time and error. At the algorithm layer, where the goal is to develop automatic methods to instantiate particular idioms, success is often judged through computational benchmarks of system time and storage. Oftentimes these unproblematic evaluation criteria can be safely treated in the objective manner typical of positivism. However, one must be alert to the emergence of cultural, political, and ethical factors that call for full adoption of an interpretivist stance. In short, a stance of epistemic pluralism is especially useful—and a variety of sensemaking models are needed—where different visualization design practices call for different theories of knowledge.
@@
When confronted with a highly political subject, such as the tendency for police data to perpetuate systemic racial violence and inequality, cognitivism tends to be a fairly toothless theoretical stance. Cases of social injustice make an especially pointed case for the limitations of cognitivism when it comes to the messy politics of human knowledge. We argue that a purely cognitive approach to schemas is insufficient for a nuanced model of sensemaking, despite the past success of this theoretical stance in domains of HCI where cognitivism has been an apt, effective, and relatively unproblematic stance, such as scientific visualization and interface design. In cases where this approach does prove problematic, an interpretivist stance offers tools for critical intervention in sensemaking processes.
@@
In calling for attention to moral dimensions of data science, the Iceberg Model is also aligned with D'Ignazio and Klein's [17] salutary call for data feminism, an approach to data that confronts power, embraces pluralism of perspectives, and surfaces the context in which data is collected and used. Far from undermining quantitative reasoning, an interpretivist theory of knowledge yields a more humane, ethical, and defensible stance toward sensemaking with data. We argue that a key dimension of the changes called for by D’Ignazio & Klein will be to account for the centrality, ubiquity, and variety of schemas at play in any sensemaking process.
@@ 7.3 The Virtue of Epistemic Humility
Recognizing these social entanglements entails adopting a fundamental stance of humility with respect to data. Facts are never neutral, nobody is truly impartial, and matters of knowledge are inevitably shaped by concepts, measurements, models, and tools of our own creation. But this observation does not mean the world is unknowable—far from it. An interpretivist viewpoint calls attention to the fact that knowledge originates in real human labor, ideas, institutions, and collective memory. This framing means knowledge must be situated in human context [97], treating a multiplicity of perspectives and interpretations as elements of a more comprehensive picture than any one viewpoint. In order for knowledge to be constructed soundly and communicated clearly, it must always be framed. And while every framework leaves something outside its edges, there is value in making the framing process explicit because this underlines the purposeful decisions that lead to one conclusion rather than another. Some biases neatly isolate what matters to us, just as other biases lead us astray.
@@
Epistemology is the study of knowledge itself, and epistemic humility acknowledges how easily human beings can fail or be misled in the pursuit of knowledge. Our interpretivist process model promotes the virtue of epistemic humility by foregrounding the messy complexity of knowledge as an omnipresent factor in sensemaking. This stance is especially needed in data-driven fields, where data's apparent 'givenness' is implied in the etymology of the very word 'data', Latin for what is given. Opposition to this misnomer led Drucker to suggest the neologism capta to emphasize that this information is never just given but rather captured, gathered, ordered, and given meaning through people's actions [98]. Whether we adopt the word capta or merely reform our assumptions about the nature of data, it is incumbent on all researchers to heed this distinction, particularly in a field so human-centered as visualization is.
@@
While cognitivism has proved both popular and useful at the intersection of psychology and computer science, as a broadly positivist stance it risks drifting into a reductive view of data as objective facts, transportable across contexts, carrying no perspective of their own. The philosopher Thomas Nagel [99] has critiqued this naive realist stance as the fallacy of a view from nowhere, arguing that “the subjectivity of consciousness is an irreducible feature of reality—without which we couldn't do physics or anything else.” Haraway [72] labels the same problem as the god trick, in which the visual presentation of scientific data gives the impression of an all-encompassing understanding like that of an omniscient deity. When data-driven inquiry goes astray, the cause is often this pursuit of an objective view from nowhere without attending to subjective and socially embedded dimensions of data like its context, construction, and ultimate limitations.
@@
Although positivistic research methodologies remain sensible approaches to many scientific subjects, we urge visualization researchers to recognize that data is an artifact that ultimately emerges from human interpretation and is often enmeshed in society and politics. Many visualization researchers already do, at least tacitly. To reserve this epistemic humility fosters dynamic adaptability in cases where positivism clearly fails us. The ideal outcome is to continue developing the powerful affordances of visualization while embracing a supple understanding of data as contingent, constructed, and contestable.
@@ 8 Conclusion
Data analysis offers undeniably powerful tools for visualization research, but the political complexities of human knowledge present ineluctable, though not intractable, problems. For all the utility and appeal of data as the basis for so much scientific inquiry, the prevailing conception of data as given should challenge us to give closer attention to the social impact of data-driven reasoning and humanistic alternatives at hand. Data begins with human sensemaking even when its ultimate impact increasingly emerges from automated systems using algorithms, ML, and AI. Matters of race, gender, policing, and economic inequality pose especially difficult challenges, but there is real promise in acknowledging the fundamental role of interpretation and the complicating effects of power in the sensemaking process. We thus offer an interpretivist model of sensemaking where datasets grow out of the schemas that give them structure, meaning, and context, not vice versa. We distinguish between explicit and tacit schemas, rejecting the treatment of schemas as monolithic structures that previous models place late in the sensemaking process. We also emphasize the importance of examining multiple schemas, to avoid the limitations of a single perspective. We showcase the Iceberg Model's descriptive and prescriptive capacity through four scenarios that map the successes and failures of the sensemaking process to the three phases and nine actions of the model. The result is an interpretivist framework that incorporates anti-positivist criticism of data-centric reasoning without rejecting data analysis altogether.
Acknowledgment
The authors wish to thank Miriah Meyer, Daniel Weiskopf, Ben Shneiderman, Irina Shklovski, Mark Hansen, Nick Diakopoulos, the UBC InfoVis group, and the Vis Collective at Linköping University for helpful feedback. Steve Kasica contributed to an early iteration of this model that focused on the work of data journalists. Rosalie Yu provided additional design work for Figure 4. This work was supported in part by a grant from the Wallenberg AI and Autonomous Systems Program (WASP) and by NSERC DG RGPIN-2014-06309.

### v4

@@ 7.1 Limitations
A fundamental limitation of all interpretivist approaches, including our own, is that centering subjectivity necessarily limits the reproducibility and replicability of results [20]. Two analysts who use Iceberg Sensemaking may well come to different conclusions. Our model locates this difference in the acknowledgement of tacit schemas, where the specific information one chooses to articulate from the mass of submerged tacit schemas may be entirely contingent, leading to different sensemaking outcomes between different analysts in varying conditions and contexts. When reproducible results are required, a broadly positivist approach is still tenable under this model because the epistemic pluralism of Iceberg Sensemaking is inclusive of positivism, as well—just as long as the limitations of a positivist approach have been sufficiently scrutinized through the broader interpretivist framework of this model. Moreover, positivist approaches will typically be more time-efficient than the slower interpretivist approaches that demand scrutiny of the received wisdom and prevailing expectations of a scientific field.
@@ 7.2 The Value of an Interpretivist Model
In building this model, we have made an active effort to describe how positivist approaches to data analysis work when they work, while also explaining how they go wrong. A sensemaking model should be just as good at diagnosing how the process fails as describing how it succeeds. While the positivism of past sensemaking models is often implicit, it is most evident in the absence of consideration for faulty and misleading data. Correcting this limitation requires more than just the inclusion of a cleaning phase, because the deepest problems with many datasets only emerge through scrutiny of conceptual frameworks that are taken for granted.
@@
As a rule of thumb, the more political a subject, the more fraught the implications of tacit schemas in data analysis, and the greater the need for interpretivist approaches to sensemaking. For example, court records may seem like straightforward data, but they carry a tacit schema reflecting the entire structure of the criminal justice system. Likewise, calibrating a sensor to measure water quality may seem straightforward, but making use of the readings to assess public health concerns involves acts of interpretation that are deeply embedded in social context. While data entry and automated data gathering could be seen as passive tasks that simply require application of the intended cognitive schema, sensemaking with any dataset requires active attention to the revision of explicit schemas through articulation of tacit schemas. Although biased data is typically discussed in the context of ethics, it is just as much an epistemic matter when unfairness is rooted in failures of investigative rigor.
@@
In the specific case of visualization, the Nested Model of visualization design [103] can be a useful guide for when to take an interpretivist or cognitivist stance. In the Nested Model, the abstraction layer is where the designer must make judgements about the tasks and data of target users, and this is prime territory for an interpretivist approach. Problem-driven design and requirements elicitation are processes that necessarily involve interpretation, judgment, and subjectivity. In contrast, the lower two levels of the Nested Model are more amenable to a cognitivist approach. At the idiom layer, which deals with visual encoding and interaction design choices, success is often gauged using methods such as controlled laboratory experiments where human performance is measured in terms of time and error. At the algorithm layer, where the goal is to develop automatic methods to instantiate particular idioms, success is often judged through computational benchmarks of system time and storage. Oftentimes these unproblematic evaluation criteria can be safely treated in the objective manner typical of positivism. However, one must be alert to the emergence of cultural, political, and ethical factors that call for full adoption of an interpretivist stance. In short, a stance of epistemic pluralism is especially useful—and a variety of sensemaking models are needed—where different visualization design practices call for different theories of knowledge.
@@
When confronted with a highly political subject, such as the tendency for police data to perpetuate systemic racial violence and inequality, cognitivism tends to be a fairly toothless theoretical stance. Cases of social injustice make an especially pointed case for the limitations of cognitivism when it comes to the messy politics of human knowledge. We argue that a purely cognitive approach to schemas is insufficient for a nuanced model of sensemaking, despite the past success of this theoretical stance in domains of visualization and HCI where cognitivism has been an apt, effective, and relatively unproblematic stance, such as scientific visualization and interface design. In cases where a cognitivist or broadly positivist approach does prove problematic, an interpretivist stance offers tools for critical intervention in the sensemaking process.
@@
In calling for attention to ethical dimensions of data science, the Iceberg Model is also aligned with D'Ignazio and Klein’s [17] salutary call for data feminism, an approach to data that confronts power, embraces pluralism of perspectives, and surfaces the context in which data is collected and used. Far from undermining quantitative reasoning, an interpretivist theory of knowledge yields a more humane, ethical, and defensible stance toward sensemaking with data. We argue that a key dimension of the changes called for by D’Ignazio & Klein will be to account for the centrality, ubiquity, and variety of schemas at play in any sensemaking process.
@@ 7.3 The Virtue of Epistemic Humility
Recognizing the social entanglements of data entails adopting a fundamental stance of humility with respect to data-driven reasoning. Facts are never neutral, nobody is truly impartial, and matters of knowledge are inevitably shaped by concepts, measurements, models, and tools of our own creation. But this observation does not mean the world is unknowable—far from it. An interpretivist viewpoint calls attention to the fact that knowledge originates in real human labor, ideas, institutions, and collective memory. This framing implies knowledge must be situated in human context, treating a multiplicity of perspectives and interpretations as elements of a more comprehensive picture than any one viewpoint [104]. In order for knowledge to be constructed soundly and communicated clearly, it must always be carefully framed. And while every framework leaves something outside its edges, there is value in making the framing process explicit because this underlines the purposeful decisions that lead to one conclusion rather than another. Some biases neatly isolate what matters to us, just as other biases lead us astray.
@@
Epistemology is the study of knowledge itself, and epistemic humility acknowledges how easily human beings can fail or be misled in the pursuit of knowledge. Our interpretivist process model promotes the virtue of epistemic humility by foregrounding the messy complexity of knowledge as an ever-present factor in sensemaking. This stance is especially needed in data-driven fields, where data’s apparent givenness is implied in the etymology of the very word ‘data’, which is Latin for ‘what is given’. Opposition to this misnomer led Drucker to suggest the neologism capta to emphasize that information is never just given but rather captured, gathered, ordered, and given meaning through people’s actions [105]. Whether researchers adopt the word ‘capta’ or merely reform our assumptions about the nature of data, it is incumbent on all researchers to heed this distinction, particularly in a field so human-centered as visualization is.
@@
While cognitivism has proved both popular and useful at the intersection of psychology and computer science, as a broadly positivist stance it risks drifting into a reductive view of data as objective facts, transportable across contexts, carrying no perspective of their own. The philosopher Nagel [106] has critiqued this naive realist stance as the fallacy of a view from nowhere, arguing that “the subjectivity of consciousness is an irreducible feature of reality—without which we couldn’t do physics or anything else”. Haraway [73] labels the same problem as the god trick, in which the visual presentation of scientific data gives the impression of an all-encompassing understanding like that of an omniscient deity. When data-driven inquiry goes astray, the cause is often this pursuit of an objective view from nowhere without attending to subjective and socially embedded dimensions of data like its context, construction, and ultimate limitations.
@@
Although positivist research methodologies remain sensible approaches to many scientific subjects, we urge visualization researchers to recognize that this stance is just one option. Taking a different stance of epistemic humility, data is an artifact that ultimately emerges from human interpretation and is often enmeshed in society and politics. Many visualization researchers already take this stance to some extent, valuing contextual and interpretive factors even in studies with the outward appearance of positivist neutrality. To actively promote epistemic humility would mean letting go of that apparent neutrality and instead taking a pluralist stance toward visualization studies. The ideal outcome will be to continue developing the powerful affordances of visualization while embracing a supple understanding of data as contingent, constructed, and contestable.
@@ 8 Conclusion
Data analysis offers undeniably powerful tools for visualization research, but the political complexities of human knowledge present ineluctable, though not intractable, problems. For all the utility and appeal of data as the basis for so much scientific inquiry, the prevailing conception of data as given should challenge us to give closer attention to the social impact of data-driven reasoning and humanistic alternatives at hand. Data begins with human sensemaking even when its ultimate impact increasingly emerges from automated systems using algorithms, ML, and AI. Matters of race, gender, policing, and economic inequality pose especially difficult challenges, but there is real promise in acknowledging the fundamental role of human interpretation and the complicating effects of power in the sensemaking process. We thus offer an interpretivist model of sensemaking where datasets grow out of the schemas that give them structure, meaning, and context, not vice versa. We distinguish between explicit and tacit schemas, rejecting the treatment of schemas as monolithic structures that previous models place late in the sensemaking process. We also emphasize the importance of examining multiple schemas to avoid the limitations of a single perspective. We showcase the Iceberg Model’s descriptive and prescriptive capacity through four scenarios that map the successes and failures of the sensemaking process to the three phases and nine actions of the model. The result is an interpretivist framework that mobilizes anti-positivist critiques of data-centric reasoning without rejecting data analysis altogether.
Acknowledgments
The authors wish to thank Miriah Meyer, Daniel Weiskopf, Ben Shneiderman, Irina Shklovski, Mark Hansen, Nick Diakopoulos, the UBC InfoVis group, the Vis Collective at Linköping University, and the anonymous reviewers of this paper for their helpful feedback. Steve Kasica contributed to an early iteration of this model that focused on the work of data journalists. Rosalie Yu provided additional design work for Figure 4. This work was supported in part by a grant from the Wallenberg AI and Autonomous Systems Program (WASP) and by NSERC DG RGPIN-2014-06309.